### v1

# Measurement of the branching fraction of $\eta\rightarrow\mu^{+}\mu^{-}$ and search for $\eta\rightarrow e^{+}e^{-}$

Measurement of the branching fraction of $\eta\rightarrow\mu^{+}\mu^{-}$ and search for $\eta\rightarrow e^{+}e^{-}$
M. Ablikim1[Uncaptioned image], M. N. Achasov4,b[Uncaptioned image], P. Adlarson77[Uncaptioned image], X. C. Ai82[Uncaptioned image], R. Aliberti36[Uncaptioned image], A. Amoroso76A,76C[Uncaptioned image], Q. An73,59,†, Y. Bai58[Uncaptioned image], O. Bakina37[Uncaptioned image], Y. Ban47,g[Uncaptioned image], H.-R. Bao65[Uncaptioned image], V. Batozskaya1,45[Uncaptioned image], K. Begzsuren33, N. Berger36[Uncaptioned image], M. Berlowski45[Uncaptioned image], M. Bertani29A[Uncaptioned image], D. Bettoni30A[Uncaptioned image], F. Bianchi76A,76C[Uncaptioned image], E. Bianco76A,76C, A. Bortone76A,76C[Uncaptioned image], I. Boyko37[Uncaptioned image], R. A. Briere5[Uncaptioned image], A. Brueggemann70[Uncaptioned image], H. Cai78[Uncaptioned image], M. H. Cai39,j,k[Uncaptioned image], X. Cai1,59[Uncaptioned image], A. Calcaterra29A[Uncaptioned image], G. F. Cao1,65[Uncaptioned image], N. Cao1,65[Uncaptioned image], S. A. Cetin63A[Uncaptioned image], X. Y. Chai47,g[Uncaptioned image], J. F. Chang1,59[Uncaptioned image], G. R. Che44[Uncaptioned image], Y. Z. Che1,59,65[Uncaptioned image], C. H. Chen9[Uncaptioned image], Chao Chen56[Uncaptioned image], G. Chen1[Uncaptioned image], H. S. Chen1,65[Uncaptioned image], H. Y. Chen21[Uncaptioned image], M. L. Chen1,59,65[Uncaptioned image], S. J. Chen43[Uncaptioned image], S. L. Chen46[Uncaptioned image], S. M. Chen62[Uncaptioned image], T. Chen1,65[Uncaptioned image], X. R. Chen32,65[Uncaptioned image], X. T. Chen1,65[Uncaptioned image], X. Y. Chen12,f[Uncaptioned image], Y. B. Chen1,59[Uncaptioned image], Y. Q. Chen35[Uncaptioned image], Y. Q. Chen16[Uncaptioned image], Z. Chen25[Uncaptioned image], Z. J. Chen26,h[Uncaptioned image], Z. K. Chen60[Uncaptioned image], J. C. Cheng46[Uncaptioned image], S. K. Choi10[Uncaptioned image], X. Chu12,f[Uncaptioned image], G. Cibinetto30A[Uncaptioned image], F. Cossio76C[Uncaptioned image], J. Cottee-Meldrum64[Uncaptioned image], J. J. Cui51[Uncaptioned image], H. L. Dai1,59[Uncaptioned image], J. P. Dai80[Uncaptioned image], A. Dbeyssi19, R. E. de Boer3[Uncaptioned image], D. Dedovich37[Uncaptioned image], C. Q. Deng74[Uncaptioned image], Z. Y. Deng1[Uncaptioned image], A. Denig36[Uncaptioned image], I. Denysenko37[Uncaptioned image], M. Destefanis76A,76C[Uncaptioned image], F. De Mori76A,76C[Uncaptioned image], B. Ding68,1[Uncaptioned image], X. X. Ding47,g[Uncaptioned image], Y. Ding41[Uncaptioned image], Y. Ding35[Uncaptioned image], Y. X. Ding31[Uncaptioned image], J. Dong1,59[Uncaptioned image], L. Y. Dong1,65[Uncaptioned image], M. Y. Dong1,59,65[Uncaptioned image], X. Dong78[Uncaptioned image], M. C. Du1[Uncaptioned image], S. X. Du82[Uncaptioned image], S. X. Du12,f[Uncaptioned image], Y. Y. Duan56[Uncaptioned image], Z. H. Duan43[Uncaptioned image], P. Egorov37,a[Uncaptioned image], G. F. Fan43[Uncaptioned image], J. J. Fan20[Uncaptioned image], Y. H. Fan46[Uncaptioned image], J. Fang1,59[Uncaptioned image], J. Fang60[Uncaptioned image], S. S. Fang1,65[Uncaptioned image], W. X. Fang1[Uncaptioned image], Y. Q. Fang1,59[Uncaptioned image], L. Fava76B,76C[Uncaptioned image], F. Feldbauer3[Uncaptioned image], G. Felici29A[Uncaptioned image], C. Q. Feng73,59[Uncaptioned image], J. H. Feng16[Uncaptioned image], L. Feng39,j,k[Uncaptioned image], Q. X. Feng39,j,k[Uncaptioned image], Y. T. Feng73,59[Uncaptioned image], M. Fritsch3[Uncaptioned image], C. D. Fu1[Uncaptioned image], J. L. Fu65[Uncaptioned image], Y. W. Fu1,65[Uncaptioned image], H. Gao65[Uncaptioned image], X. B. Gao42[Uncaptioned image], Y. Gao73,59[Uncaptioned image], Y. N. Gao47,g[Uncaptioned image], Y. N. Gao20[Uncaptioned image], Y. Y. Gao31[Uncaptioned image], S. Garbolino76C[Uncaptioned image], I. Garzia30A,30B[Uncaptioned image], L. Ge58[Uncaptioned image], P. T. Ge20[Uncaptioned image], Z. W. Ge43[Uncaptioned image], C. Geng60[Uncaptioned image], E. M. Gersabeck69[Uncaptioned image], A. Gilman71[Uncaptioned image], K. Goetzen13[Uncaptioned image], J. D. Gong35[Uncaptioned image], L. Gong41[Uncaptioned image], W. X. Gong1,59[Uncaptioned image], W. Gradl36[Uncaptioned image], S. Gramigna30A,30B[Uncaptioned image], M. Greco76A,76C[Uncaptioned image], M. H. Gu1,59[Uncaptioned image], Y. T. Gu15[Uncaptioned image], C. Y. Guan1,65[Uncaptioned image], A. Q. Guo32[Uncaptioned image], L. B. Guo42[Uncaptioned image], M. J. Guo51[Uncaptioned image], R. P. Guo50[Uncaptioned image], Y. P. Guo12,f[Uncaptioned image], A. Guskov37,a[Uncaptioned image], J. Gutierrez28[Uncaptioned image], K. L. Han65[Uncaptioned image], T. T. Han1[Uncaptioned image], F. Hanisch3[Uncaptioned image], K. D. Hao73,59[Uncaptioned image], X. Q. Hao20[Uncaptioned image], F. A. Harris67[Uncaptioned image], K. K. He56[Uncaptioned image], K. L. He1,65[Uncaptioned image], F. H. Heinsius3[Uncaptioned image], C. H. Heinz36[Uncaptioned image], Y. K. Heng1,59,65[Uncaptioned image], C. Herold61[Uncaptioned image], P. C. Hong35[Uncaptioned image], G. Y. Hou1,65[Uncaptioned image], X. T. Hou1,65[Uncaptioned image], Y. R. Hou65[Uncaptioned image], Z. L. Hou1[Uncaptioned image], H. M. Hu1,65[Uncaptioned image], J. F. Hu57,i[Uncaptioned image], Q. P. Hu73,59[Uncaptioned image], S. L. Hu12,f[Uncaptioned image], T. Hu1,59,65[Uncaptioned image], Y. Hu1[Uncaptioned image], Z. M. Hu60[Uncaptioned image], G. S. Huang73,59[Uncaptioned image], K. X. Huang60[Uncaptioned image], L. Q. Huang32,65[Uncaptioned image], P. Huang43[Uncaptioned image], X. T. Huang51[Uncaptioned image], Y. P. Huang1[Uncaptioned image], Y. S. Huang60[Uncaptioned image], T. Hussain75[Uncaptioned image], N. Hüsken36[Uncaptioned image], N. in der Wiesche70[Uncaptioned image], J. Jackson28[Uncaptioned image], Q. Ji1[Uncaptioned image], Q. P. Ji20[Uncaptioned image], W. Ji1,65[Uncaptioned image], X. B. Ji1,65[Uncaptioned image], X. L. Ji1,59[Uncaptioned image], Y. Y. Ji51[Uncaptioned image], Z. K. Jia73,59[Uncaptioned image], D. Jiang1,65[Uncaptioned image], H. B. Jiang78[Uncaptioned image], P. C. Jiang47,g[Uncaptioned image], S. J. Jiang9[Uncaptioned image], T. J. Jiang17[Uncaptioned image], X. S. Jiang1,59,65[Uncaptioned image], Y. Jiang65[Uncaptioned image], J. B. Jiao51[Uncaptioned image], J. K. Jiao35[Uncaptioned image], Z. Jiao24[Uncaptioned image], S. Jin43[Uncaptioned image], Y. Jin68[Uncaptioned image], M. Q. Jing1,65[Uncaptioned image], X. M. Jing65[Uncaptioned image], T. Johansson77[Uncaptioned image], S. Kabana34[Uncaptioned image], N. Kalantar-Nayestanaki66[Uncaptioned image], X. L. Kang9[Uncaptioned image], X. S. Kang41[Uncaptioned image], M. Kavatsyuk66[Uncaptioned image], B. C. Ke82[Uncaptioned image], V. Khachatryan28[Uncaptioned image], A. Khoukaz70[Uncaptioned image], R. Kiuchi1, O. B. Kolcu63A[Uncaptioned image], B. Kopf3[Uncaptioned image], M. Kuessner3[Uncaptioned image], X. Kui1,65[Uncaptioned image], N. Kumar27[Uncaptioned image], A. Kupsc45,77[Uncaptioned image], W. Kühn38[Uncaptioned image], Q. Lan74[Uncaptioned image], W. N. Lan20[Uncaptioned image], T. T. Lei73,59[Uncaptioned image], M. Lellmann36[Uncaptioned image], T. Lenz36[Uncaptioned image], C. Li48[Uncaptioned image], C. Li44[Uncaptioned image], C. H. Li40[Uncaptioned image], C. K. Li21[Uncaptioned image], D. M. Li82[Uncaptioned image], F. Li1,59[Uncaptioned image], G. Li1[Uncaptioned image], H. B. Li1,65[Uncaptioned image], H. J. Li20[Uncaptioned image], H. N. Li57,i[Uncaptioned image], Hui Li44[Uncaptioned image], J. R. Li62[Uncaptioned image], J. S. Li60[Uncaptioned image], K. Li1[Uncaptioned image], K. L. Li20[Uncaptioned image], K. L. Li39,j,k[Uncaptioned image], L. J. Li1,65[Uncaptioned image], Lei Li49[Uncaptioned image], M. H. Li44[Uncaptioned image], M. R. Li1,65[Uncaptioned image], P. L. Li65[Uncaptioned image], P. R. Li39,j,k[Uncaptioned image], Q. M. Li1,65[Uncaptioned image], Q. X. Li51[Uncaptioned image], R. Li18,32[Uncaptioned image], S. X. Li12[Uncaptioned image], T. Li51[Uncaptioned image], T. Y. Li44[Uncaptioned image], W. D. Li1,65[Uncaptioned image], W. G. Li1,†[Uncaptioned image], X. Li1,65[Uncaptioned image], X. H. Li73,59[Uncaptioned image], X. L. Li51[Uncaptioned image], X. Y. Li1,8[Uncaptioned image], X. Z. Li60[Uncaptioned image], Y. Li20[Uncaptioned image], Y. G. Li47,g[Uncaptioned image], Y. P. Li35[Uncaptioned image], Z. J. Li60[Uncaptioned image], Z. Y. Li80[Uncaptioned image], C. Liang43[Uncaptioned image], H. Liang73,59[Uncaptioned image], Y. F. Liang55[Uncaptioned image], Y. T. Liang32,65[Uncaptioned image], G. R. Liao14[Uncaptioned image], L. B. Liao60[Uncaptioned image], M. H. Liao60[Uncaptioned image], Y. P. Liao1,65[Uncaptioned image], J. Libby27[Uncaptioned image], A. Limphirat61[Uncaptioned image], C. C. Lin56[Uncaptioned image], D. X. Lin32,65[Uncaptioned image], L. Q. Lin40[Uncaptioned image], T. Lin1[Uncaptioned image], B. J. Liu1[Uncaptioned image], B. X. Liu78[Uncaptioned image], C. Liu35[Uncaptioned image], C. X. Liu1[Uncaptioned image], F. Liu1[Uncaptioned image], F. H. Liu54[Uncaptioned image], Feng Liu6[Uncaptioned image], G. M. Liu57,i[Uncaptioned image], H. Liu39,j,k[Uncaptioned image], H. B. Liu15[Uncaptioned image], H. H. Liu1[Uncaptioned image], H. M. Liu1,65[Uncaptioned image], Huihui Liu22[Uncaptioned image], J. B. Liu73,59[Uncaptioned image], J. J. Liu21[Uncaptioned image], K. Liu39,j,k[Uncaptioned image], K. Liu74[Uncaptioned image], K. Y. Liu41[Uncaptioned image], Ke Liu23[Uncaptioned image], L. C. Liu44[Uncaptioned image], Lu Liu44[Uncaptioned image], M. H. Liu12,f[Uncaptioned image], M. H. Liu35[Uncaptioned image], P. L. Liu1[Uncaptioned image], Q. Liu65[Uncaptioned image], S. B. Liu73,59[Uncaptioned image], T. Liu12,f[Uncaptioned image], W. K. Liu44[Uncaptioned image], W. M. Liu73,59[Uncaptioned image], W. T. Liu40[Uncaptioned image], X. Liu39,j,k[Uncaptioned image], X. Liu40[Uncaptioned image], X. K. Liu39,j,k[Uncaptioned image], X. L. Liu12,f[Uncaptioned image], X. Y. Liu78[Uncaptioned image], Y. Liu39,j,k[Uncaptioned image], Y. Liu82[Uncaptioned image], Yuan Liu82[Uncaptioned image], Y. B. Liu44[Uncaptioned image], Z. A. Liu1,59,65[Uncaptioned image], Z. D. Liu9[Uncaptioned image], Z. Q. Liu51[Uncaptioned image], X. C. Lou1,59,65[Uncaptioned image], F. X. Lu60[Uncaptioned image], H. J. Lu24[Uncaptioned image], J. G. Lu1,59[Uncaptioned image], X. L. Lu16[Uncaptioned image], Y. Lu7[Uncaptioned image], Y. H. Lu1,65[Uncaptioned image], Y. P. Lu1,59[Uncaptioned image], Z. H. Lu1,65[Uncaptioned image], C. L. Luo42[Uncaptioned image], J. R. Luo60[Uncaptioned image], J. S. Luo1,65[Uncaptioned image], M. X. Luo81, T. Luo12,f[Uncaptioned image], X. L. Luo1,59[Uncaptioned image], Z. Y. Lv23[Uncaptioned image], X. R. Lyu65,o[Uncaptioned image], Y. F. Lyu44[Uncaptioned image], Y. H. Lyu82[Uncaptioned image], F. C. Ma41[Uncaptioned image], H. L. Ma1[Uncaptioned image], Heng Ma26,h[Uncaptioned image], J. L. Ma1,65[Uncaptioned image], L. L. Ma51[Uncaptioned image], L. R. Ma68[Uncaptioned image], Q. M. Ma1[Uncaptioned image], R. Q. Ma1,65[Uncaptioned image], R. Y. Ma20[Uncaptioned image], T. Ma73,59[Uncaptioned image], X. T. Ma1,65[Uncaptioned image], X. Y. Ma1,59[Uncaptioned image], Y. M. Ma32[Uncaptioned image], F. E. Maas19[Uncaptioned image], I. MacKay71[Uncaptioned image], M. Maggiora76A,76C[Uncaptioned image], S. Malde71[Uncaptioned image], Q. A. Malik75[Uncaptioned image], H. X. Mao39,j,k[Uncaptioned image], Y. J. Mao47,g[Uncaptioned image], Z. P. Mao1[Uncaptioned image], S. Marcello76A,76C[Uncaptioned image], A. Marshall64[Uncaptioned image], F. M. Melendi30A,30B[Uncaptioned image], Y. H. Meng65[Uncaptioned image], Z. X. Meng68[Uncaptioned image], G. Mezzadri30A[Uncaptioned image], H. Miao1,65[Uncaptioned image], T. J. Min43[Uncaptioned image], R. E. Mitchell28[Uncaptioned image], X. H. Mo1,59,65[Uncaptioned image], B. Moses28[Uncaptioned image], N. Yu. Muchnoi4,b[Uncaptioned image], J. Muskalla36[Uncaptioned image], Y. Nefedov37[Uncaptioned image], F. Nerling19,d[Uncaptioned image], L. S. Nie21[Uncaptioned image], I. B. Nikolaev4,b, Z. Ning1,59[Uncaptioned image], S. Nisar11,l, Q. L. Niu39,j,k[Uncaptioned image], W. D. Niu12,f[Uncaptioned image], C. Normand64[Uncaptioned image], S. L. Olsen10,65[Uncaptioned image], Q. Ouyang1,59,65[Uncaptioned image], S. Pacetti29B,29C[Uncaptioned image], X. Pan56[Uncaptioned image], Y. Pan58[Uncaptioned image], A. Pathak10[Uncaptioned image], Y. P. Pei73,59[Uncaptioned image], M. Pelizaeus3[Uncaptioned image], H. P. Peng73,59[Uncaptioned image], X. J. Peng39,j,k[Uncaptioned image], Y. Y. Peng39,j,k[Uncaptioned image], K. Peters13,d[Uncaptioned image], K. Petridis64[Uncaptioned image], J. L. Ping42[Uncaptioned image], R. G. Ping1,65[Uncaptioned image], S. Plura36[Uncaptioned image], V. Prasad35[Uncaptioned image], F. Z. Qi1[Uncaptioned image], H. R. Qi62[Uncaptioned image], M. Qi43[Uncaptioned image], S. Qian1,59[Uncaptioned image], W. B. Qian65[Uncaptioned image], C. F. Qiao65[Uncaptioned image], J. H. Qiao20[Uncaptioned image], J. J. Qin74[Uncaptioned image], J. L. Qin56[Uncaptioned image], L. Q. Qin14[Uncaptioned image], L. Y. Qin73,59[Uncaptioned image], P. B. Qin74[Uncaptioned image], X. P. Qin12,f[Uncaptioned image], X. S. Qin51[Uncaptioned image], Z. H. Qin1,59[Uncaptioned image], J. F. Qiu1[Uncaptioned image], Z. H. Qu74[Uncaptioned image], J. Rademacker64[Uncaptioned image], C. F. Redmer36[Uncaptioned image], A. Rivetti76C[Uncaptioned image], M. Rolo76C[Uncaptioned image], G. Rong1,65[Uncaptioned image], S. S. Rong1,65[Uncaptioned image], F. Rosini29B,29C[Uncaptioned image], Ch. Rosner19[Uncaptioned image], M. Q. Ruan1,59[Uncaptioned image], N. Salone45,p[Uncaptioned image], A. Sarantsev37,c[Uncaptioned image], Y. Schelhaas36[Uncaptioned image], K. Schoenning77[Uncaptioned image], M. Scodeggio30A[Uncaptioned image], K. Y. Shan12,f[Uncaptioned image], W. Shan25[Uncaptioned image], X. Y. Shan73,59[Uncaptioned image], Z. J. Shang39,j,k[Uncaptioned image], J. F. Shangguan17[Uncaptioned image], L. G. Shao1,65[Uncaptioned image], M. Shao73,59[Uncaptioned image], C. P. Shen12,f[Uncaptioned image], H. F. Shen1,8[Uncaptioned image], W. H. Shen65[Uncaptioned image], X. Y. Shen1,65[Uncaptioned image], B. A. Shi65[Uncaptioned image], H. Shi73,59[Uncaptioned image], J. L. Shi12,f[Uncaptioned image], J. Y. Shi1[Uncaptioned image], S. Y. Shi74[Uncaptioned image], X. Shi1,59[Uncaptioned image], H. L. Song73,59[Uncaptioned image], J. J. Song20[Uncaptioned image], T. Z. Song60[Uncaptioned image], W. M. Song35[Uncaptioned image], Y. J. Song12,f[Uncaptioned image], Y. X. Song47,g,m[Uncaptioned image], Zirong Song26,h[Uncaptioned image], S. Sosio76A,76C[Uncaptioned image], S. Spataro76A,76C[Uncaptioned image], S. Stansilaus71[Uncaptioned image], F. Stieler36[Uncaptioned image], S. S Su41[Uncaptioned image], Y. J. Su65[Uncaptioned image], G. B. Sun78[Uncaptioned image], G. X. Sun1[Uncaptioned image], H. Sun65[Uncaptioned image], H. K. Sun1[Uncaptioned image], J. F. Sun20[Uncaptioned image], K. Sun62[Uncaptioned image], L. Sun78[Uncaptioned image], S. S. Sun1,65[Uncaptioned image], T. Sun52,e[Uncaptioned image], Y. C. Sun78[Uncaptioned image], Y. H. Sun31[Uncaptioned image], Y. J. Sun73,59[Uncaptioned image], Y. Z. Sun1[Uncaptioned image], Z. Q. Sun1,65[Uncaptioned image], Z. T. Sun51[Uncaptioned image], C. J. Tang55, G. Y. Tang1[Uncaptioned image], J. Tang60[Uncaptioned image], J. J. Tang73,59[Uncaptioned image], L. F. Tang40[Uncaptioned image], Y. A. Tang78[Uncaptioned image], L. Y. Tao74[Uncaptioned image], M. Tat71[Uncaptioned image], J. X. Teng73,59[Uncaptioned image], J. Y. Tian73,59[Uncaptioned image], W. H. Tian60[Uncaptioned image], Y. Tian32[Uncaptioned image], Z. F. Tian78[Uncaptioned image], I. Uman63B[Uncaptioned image], B. Wang1[Uncaptioned image], B. Wang60[Uncaptioned image], Bo Wang73,59[Uncaptioned image], C. Wang39,j,k[Uncaptioned image], C. Wang20[Uncaptioned image], Cong Wang23[Uncaptioned image], D. Y. Wang47,g[Uncaptioned image], H. J. Wang39,j,k[Uncaptioned image], J. J. Wang78[Uncaptioned image], K. Wang1,59[Uncaptioned image], L. L. Wang1[Uncaptioned image], L. W. Wang35[Uncaptioned image], M. Wang51[Uncaptioned image], M. Wang73,59[Uncaptioned image], N. Y. Wang65[Uncaptioned image], S. Wang12,f[Uncaptioned image], T. Wang12,f[Uncaptioned image], T. J. Wang44[Uncaptioned image], W. Wang60[Uncaptioned image], Wei Wang74[Uncaptioned image], W. P. Wang36[Uncaptioned image], X. Wang47,g[Uncaptioned image], X. F. Wang39,j,k[Uncaptioned image], X. J. Wang40[Uncaptioned image], X. L. Wang12,f[Uncaptioned image], X. N. Wang1,65[Uncaptioned image], Y. Wang62[Uncaptioned image], Y. D. Wang46[Uncaptioned image], Y. F. Wang1,8,65[Uncaptioned image], Y. H. Wang39,j,k[Uncaptioned image], Y. J. Wang73,59[Uncaptioned image], Y. L. Wang20[Uncaptioned image], Y. N. Wang78[Uncaptioned image], Y. Q. Wang1[Uncaptioned image], Yaqian Wang18[Uncaptioned image], Yi Wang62[Uncaptioned image], Yuan Wang18,32[Uncaptioned image], Z. Wang1,59[Uncaptioned image], Z. L. Wang74[Uncaptioned image], Z. L. Wang2[Uncaptioned image], Z. Q. Wang12,f[Uncaptioned image], Z. Y. Wang1,65[Uncaptioned image], D. H. Wei14[Uncaptioned image], H. R. Wei44[Uncaptioned image], F. Weidner70[Uncaptioned image], S. P. Wen1[Uncaptioned image], Y. R. Wen40[Uncaptioned image], U. Wiedner3[Uncaptioned image], G. Wilkinson71[Uncaptioned image], M. Wolke77, C. Wu40[Uncaptioned image], J. F. Wu1,8[Uncaptioned image], L. H. Wu1[Uncaptioned image], L. J. Wu1,65[Uncaptioned image], L. J. Wu20[Uncaptioned image], Lianjie Wu20[Uncaptioned image], S. G. Wu1,65[Uncaptioned image], S. M. Wu65[Uncaptioned image], X. Wu12,f[Uncaptioned image], X. H. Wu35[Uncaptioned image], Y. J. Wu32[Uncaptioned image], Z. Wu1,59[Uncaptioned image], L. Xia73,59[Uncaptioned image], X. M. Xian40[Uncaptioned image], B. H. Xiang1,65[Uncaptioned image], D. Xiao39,j,k[Uncaptioned image], G. Y. Xiao43[Uncaptioned image], H. Xiao74[Uncaptioned image], Y. L. Xiao12,f[Uncaptioned image], Z. J. Xiao42[Uncaptioned image], C. Xie43[Uncaptioned image], K. J. Xie1,65[Uncaptioned image], X. H. Xie47,g[Uncaptioned image], Y. Xie51[Uncaptioned image], Y. G. Xie1,59[Uncaptioned image], Y. H. Xie6[Uncaptioned image], Z. P. Xie73,59[Uncaptioned image], T. Y. Xing1,65[Uncaptioned image], C. F. Xu1,65, C. J. Xu60[Uncaptioned image], G. F. Xu1[Uncaptioned image], H. Y. Xu68,2[Uncaptioned image], H. Y. Xu2[Uncaptioned image], M. Xu73,59[Uncaptioned image], Q. J. Xu17[Uncaptioned image], Q. N. Xu31[Uncaptioned image], T. D. Xu74[Uncaptioned image], W. Xu1[Uncaptioned image], W. L. Xu68[Uncaptioned image], X. P. Xu56[Uncaptioned image], Y. Xu41[Uncaptioned image], Y. Xu12,f[Uncaptioned image], Y. C. Xu79[Uncaptioned image], Z. S. Xu65[Uncaptioned image], F. Yan12,f[Uncaptioned image], H. Y. Yan40[Uncaptioned image], L. Yan12,f[Uncaptioned image], W. B. Yan73,59[Uncaptioned image], W. C. Yan82[Uncaptioned image], W. H. Yan6[Uncaptioned image], W. P. Yan20[Uncaptioned image], X. Q. Yan1,65[Uncaptioned image], H. J. Yang52,e[Uncaptioned image], H. L. Yang35[Uncaptioned image], H. X. Yang1[Uncaptioned image], J. H. Yang43[Uncaptioned image], R. J. Yang20[Uncaptioned image], T. Yang1[Uncaptioned image], Y. Yang12,f[Uncaptioned image], Y. F. Yang44[Uncaptioned image], Y. H. Yang43[Uncaptioned image], Y. Q. Yang9[Uncaptioned image], Y. X. Yang1,65[Uncaptioned image], Y. Z. Yang20[Uncaptioned image], M. Ye1,59[Uncaptioned image], M. H. Ye8,†[Uncaptioned image], Z. J. Ye57,i[Uncaptioned image], Junhao Yin44[Uncaptioned image], Z. Y. You60[Uncaptioned image], B. X. Yu1,59,65[Uncaptioned image], C. X. Yu44[Uncaptioned image], G. Yu13[Uncaptioned image], J. S. Yu26,h[Uncaptioned image], L. Q. Yu12,f[Uncaptioned image], M. C. Yu41[Uncaptioned image], T. Yu74[Uncaptioned image], X. D. Yu47,g[Uncaptioned image], Y. C. Yu82[Uncaptioned image], C. Z. Yuan1,65[Uncaptioned image], H. Yuan1,65[Uncaptioned image], J. Yuan35[Uncaptioned image], J. Yuan46[Uncaptioned image], L. Yuan2[Uncaptioned image], S. C. Yuan1,65[Uncaptioned image], S. H. Yuan74[Uncaptioned image], X. Q. Yuan1[Uncaptioned image], Y. Yuan1,65[Uncaptioned image], Z. Y. Yuan60[Uncaptioned image], C. X. Yue40[Uncaptioned image], Ying Yue20[Uncaptioned image], A. A. Zafar75[Uncaptioned image], S. H. Zeng64[Uncaptioned image], X. Zeng12,f[Uncaptioned image], Y. Zeng26,h, Yujie Zeng60[Uncaptioned image], Y. J. Zeng1,65[Uncaptioned image], X. Y. Zhai35[Uncaptioned image], Y. H. Zhan60[Uncaptioned image], Shunan Zhang71[Uncaptioned image], A. Q. Zhang1,65[Uncaptioned image], B. L. Zhang1,65[Uncaptioned image], B. X. Zhang1[Uncaptioned image], D. H. Zhang44[Uncaptioned image], G. Y. Zhang20[Uncaptioned image], G. Y. Zhang1,65[Uncaptioned image], H. Zhang73,59[Uncaptioned image], H. Zhang82[Uncaptioned image], H. C. Zhang1,59,65[Uncaptioned image], H. H. Zhang60[Uncaptioned image], H. Q. Zhang1,59,65[Uncaptioned image], H. R. Zhang73,59[Uncaptioned image], H. Y. Zhang1,59[Uncaptioned image], Jin Zhang82[Uncaptioned image], J. Zhang60[Uncaptioned image], J. J. Zhang53[Uncaptioned image], J. L. Zhang21[Uncaptioned image], J. Q. Zhang42[Uncaptioned image], J. S. Zhang12,f[Uncaptioned image], J. W. Zhang1,59,65[Uncaptioned image], J. X. Zhang39,j,k[Uncaptioned image], J. Y. Zhang1[Uncaptioned image], J. Z. Zhang1,65[Uncaptioned image], Jianyu Zhang65[Uncaptioned image], L. M. Zhang62[Uncaptioned image], Lei Zhang43[Uncaptioned image], N. Zhang82[Uncaptioned image], P. Zhang1,8[Uncaptioned image], Q. Zhang20[Uncaptioned image], Q. Y. Zhang35[Uncaptioned image], R. Y. Zhang39,j,k[Uncaptioned image], S. H. Zhang1,65[Uncaptioned image], Shulei Zhang26,h[Uncaptioned image], X. M. Zhang1[Uncaptioned image], X. Y Zhang41[Uncaptioned image], X. Y. Zhang51[Uncaptioned image], Y. Zhang1[Uncaptioned image], Y. Zhang74[Uncaptioned image], Y. T. Zhang82[Uncaptioned image], Y. H. Zhang1,59[Uncaptioned image], Y. M. Zhang40[Uncaptioned image], Y. P. Zhang73,59[Uncaptioned image], Z. D. Zhang1[Uncaptioned image], Z. H. Zhang1[Uncaptioned image], Z. L. Zhang35[Uncaptioned image], Z. L. Zhang56[Uncaptioned image], Z. X. Zhang20[Uncaptioned image], Z. Y. Zhang78[Uncaptioned image], Z. Y. Zhang44[Uncaptioned image], Z. Z. Zhang46[Uncaptioned image], Zh. Zh. Zhang20[Uncaptioned image], G. Zhao1[Uncaptioned image], J. Y. Zhao1,65[Uncaptioned image], J. Z. Zhao1,59[Uncaptioned image], L. Zhao1[Uncaptioned image], L. Zhao73,59[Uncaptioned image], M. G. Zhao44[Uncaptioned image], N. Zhao80[Uncaptioned image], R. P. Zhao65[Uncaptioned image], S. J. Zhao82[Uncaptioned image], Y. B. Zhao1,59[Uncaptioned image], Y. L. Zhao56[Uncaptioned image], Y. X. Zhao32,65[Uncaptioned image], Z. G. Zhao73,59[Uncaptioned image], A. Zhemchugov37,a[Uncaptioned image], B. Zheng74[Uncaptioned image], B. M. Zheng35[Uncaptioned image], J. P. Zheng1,59[Uncaptioned image], W. J. Zheng1,65[Uncaptioned image], X. R. Zheng20[Uncaptioned image], Y. H. Zheng65,o[Uncaptioned image], B. Zhong42[Uncaptioned image], C. Zhong20[Uncaptioned image], H. Zhou36,51,n[Uncaptioned image], J. Q. Zhou35[Uncaptioned image], J. Y. Zhou35[Uncaptioned image], S. Zhou6[Uncaptioned image], X. Zhou78[Uncaptioned image], X. K. Zhou6[Uncaptioned image], X. R. Zhou73,59[Uncaptioned image], X. Y. Zhou40[Uncaptioned image], Y. X. Zhou79[Uncaptioned image], Y. Z. Zhou12,f[Uncaptioned image], A. N. Zhu65[Uncaptioned image], J. Zhu44[Uncaptioned image], K. Zhu1[Uncaptioned image], K. J. Zhu1,59,65[Uncaptioned image], K. S. Zhu12,f[Uncaptioned image], L. Zhu35[Uncaptioned image], L. X. Zhu65[Uncaptioned image], S. H. Zhu72[Uncaptioned image], T. J. Zhu12,f[Uncaptioned image], W. D. Zhu42[Uncaptioned image], W. D. Zhu12,f[Uncaptioned image], W. J. Zhu1[Uncaptioned image], W. Z. Zhu20[Uncaptioned image], Y. C. Zhu73,59[Uncaptioned image], Z. A. Zhu1,65[Uncaptioned image], X. Y. Zhuang44[Uncaptioned image], J. H. Zou1[Uncaptioned image], J. Zu73,59[Uncaptioned image] (BESIII Collaboration) 1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Guangxi University of Science and Technology, Liuzhou 545006, People’s Republic of China
17 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
18 Hebei University, Baoding 071002, People’s Republic of China
19 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
20 Henan Normal University, Xinxiang 453007, People’s Republic of China
21 Henan University, Kaifeng 475004, People’s Republic of China
22 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
23 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
24 Huangshan College, Huangshan 245000, People’s Republic of China
25 Hunan Normal University, Changsha 410081, People’s Republic of China
26 Hunan University, Changsha 410082, People’s Republic of China
27 Indian Institute of Technology Madras, Chennai 600036, India
28 Indiana University, Bloomington, Indiana 47405, USA
29 INFN Laboratori Nazionali di Frascati, (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
30 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
31 Inner Mongolia University, Hohhot 010021, People’s Republic of China
32 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
33 Institute of Physics and Technology, Mongolian Academy of Sciences, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
34 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
35 Jilin University, Changchun 130012, People’s Republic of China
36 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
37 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
38 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
39 Lanzhou University, Lanzhou 730000, People’s Republic of China
40 Liaoning Normal University, Dalian 116029, People’s Republic of China
41 Liaoning University, Shenyang 110036, People’s Republic of China
42 Nanjing Normal University, Nanjing 210023, People’s Republic of China
43 Nanjing University, Nanjing 210093, People’s Republic of China
44 Nankai University, Tianjin 300071, People’s Republic of China
45 National Centre for Nuclear Research, Warsaw 02-093, Poland
46 North China Electric Power University, Beijing 102206, People’s Republic of China
47 Peking University, Beijing 100871, People’s Republic of China
48 Qufu Normal University, Qufu 273165, People’s Republic of China
49 Renmin University of China, Beijing 100872, People’s Republic of China
50 Shandong Normal University, Jinan 250014, People’s Republic of China
51 Shandong University, Jinan 250100, People’s Republic of China
52 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
53 Shanxi Normal University, Linfen 041004, People’s Republic of China
54 Shanxi University, Taiyuan 030006, People’s Republic of China
55 Sichuan University, Chengdu 610064, People’s Republic of China
56 Soochow University, Suzhou 215006, People’s Republic of China
57 South China Normal University, Guangzhou 510006, People’s Republic of China
58 Southeast University, Nanjing 211100, People’s Republic of China
59 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
60 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
61 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
62 Tsinghua University, Beijing 100084, People’s Republic of China
63 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
64 University of Bristol, H H Wills Physics Laboratory, Tyndall Avenue, Bristol, BS8 1TL, UK
65 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
66 University of Groningen, NL-9747 AA Groningen, The Netherlands
67 University of Hawaii, Honolulu, Hawaii 96822, USA
68 University of Jinan, Jinan 250022, People’s Republic of China
69 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
70 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
71 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
72 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
73 University of Science and Technology of China, Hefei 230026, People’s Republic of China
74 University of South China, Hengyang 421001, People’s Republic of China
75 University of the Punjab, Lahore-54590, Pakistan
76 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
77 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
78 Wuhan University, Wuhan 430072, People’s Republic of China
79 Yantai University, Yantai 264005, People’s Republic of China
80 Yunnan University, Kunming 650500, People’s Republic of China
81 Zhejiang University, Hangzhou 310027, People’s Republic of China
82 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
Deceased
a Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
b Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
c Also at the NRC "Kurchatov Institute", PNPI, 188300, Gatchina, Russia
d Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
e Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
f Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
g Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
h Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
i Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
j Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
k Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
m Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
n Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
o Also at Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Hangzhou 310024, China
p Currently at Silesian University in Katowice, Chorzow, 41-500, Poland
Abstract
We analyze the decay of $\eta\rightarrow\ell^{+}\ell^{-}(\ell=e,\mu)$ via $J/\psi\rightarrow\gamma\eta^{\prime}$ and $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$, based on (10087 $\pm$ 44) $\times$ 106 $J/\psi$ events collected with the BESIII detector at the BEPCII storage rings. The branching fraction of $\eta\rightarrow\mu^{+}\mu^{-}$ is measured to be $(5.8\pm 1.0_{\rm stat}\pm 0.2_{\rm syst})\times 10^{-6}$, which is consistent with the previous measurements and theoretical expectations. In addition, no significant $\eta\to e^{+}e^{-}$ signal is observed in the $e^{+}e^{-}$ invariant mass spectrum, and an improved upper limit of ${\cal B}(\eta\to e^{+}e^{-})<2.2\times 10^{-7}$ is set at 90% confidence level.
I Introduction
Within the Standard Model (SM), the decay of $\eta$ into lepton pairs ($\eta\to\ell^{+}\ell^{-}$ with $\ell=e$ or $\mu$) proceeds via a fourth-order electromagnetic transition. The process is dominated by a two-photon intermediate state, as depicted in Fig. 1, resulting in significant suppression of the decay. Especially for $\eta\rightarrow e^{+}e^{-}$, the decay probability is expected to be very low, since the helicity factor is proportional to the square of the electron mass. Hence, a measurement of this decay probability higher than the SM model prediction could be indicative of hypothetical interactions arising from physics beyond the SM. This has stimulated considerable theoretical effort Young:1967zzb; BABU1982449; AMETLLER1983301; PhysRevD.29.911; PhysRevD.42.1509; gengLeptonicCPViolation1990; Margolis:1992gf; SAVAGE1992481; PhysRevLett.80.4633; Ametller:1993we; Knecht:1999gb; Silagadze:2006rt; Dorokhov:2009xs; Petri:2010ea devoted to calculating the decay rate of $\eta\to\ell^{+}\ell^{-}$. The predicted branching fractions for $\eta\to\mu^{+}\mu^{-}$ and $\eta\to e^{+}e^{-}$, based on the Canterbury approximants method Masjuan_2016, are $(4.72^{+0.05}_{-0.21})\times 10^{-6}$ and $(5.31^{+0.14}_{-0.04})\times 10^{-9}$, respectively.
Experimentally, a stringent upper limit of $\mathcal{B}(\eta\rightarrow e^{+}e^{-})<7\times 10^{-7}$, at 90% confidence level (CL), was obtained from the inverse process $e^{+}e^{-}\to\eta$ SND:2018egu. The world average of $\mathcal{B}(\eta\rightarrow\mu^{+}\mu^{-})=(5.7\pm 0.8)\times 10^{-6}$ PhysRevD.110.030001 is a combination of measurements performed by SATURNE II  PhysRevD.50.92 and Lepton-G DZHELYADIN1980471 more than thirty years ago. In recent years, taking advantage of the abundant production of $\eta,\eta^{\prime}$ mesons in $J/\psi$ radiative decays and the high statistics $J/\psi$ events, the BESIII experiment has made significant improvements in the study of their decays. However, the investigation of $\eta\to\ell^{+}\ell^{-}$ decays via the radiative decay $J/\psi\to\gamma\eta$ suffers from large backgrounds from $J/\psi\to\gamma\ell^{+}\ell^{-}$decays.
More recently, a novel approach Kang:2023yye was proposed to investigate $\eta$ decays, particularly for their rare or forbidden decays, with $\eta^{\prime}\rightarrow\pi\pi\eta$ produced in $J/\psi$ radiative decays. Therefore, using the sample of $(10087\pm 44)\times$ 106 BESIII:2021cxx $J/\psi$ events collected with the BESIII detector, we perform a study of $\eta\rightarrow\ell^{+}\ell^{-}$ with $J/\psi\rightarrow\gamma\eta^{\prime}$ and the subsequent decay of $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$.
Fig. 1: The Feynman diagram for $\eta\rightarrow\ell^{+}\ell^{-}$.
II BESIII DETECTOR and MONTE CARLO SIMULATION
The BESIII detector ABLIKIM2010345 is a magnetic spectrometer located at the Beijing Electron Positron Collider (BEPCII) ABLIKIM2010345. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The magnetic field was 0.9 T in 2012, which affects 11% of the total $J/\psi$ data. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identifier modules interleaved with steel. The charged-particle momentum resolution at 1 GeV/$c$ is 0.5%, and the specific ionization energy loss $\text{d}E/\text{d}x$ resolution is 6% for the electrons from Bhabha scattering. The EMC measures photon energies with a resolution of 2.5% (5%) at 1 GeV in the barrel (end-cap) region. The time resolution of the TOF barrel part is 68 ps, while that in the end-cap part is 110 ps. The end-cap TOF system was upgraded in 2015 using multi-gap resistive plate chamber technology, providing a time resolution of 60 ps Li:2017jpg; Guo:2017sjt; CAO2020163053, which benefits 87% of the data used in this paper.
Monte Carlo (MC) simulated data samples are produced with a geant4-based AGOSTINELLI2003250 software package, which includes the geometric description of the BESIII detector and the detector response. These samples are used to determine detection efficiencies and estimate the backgrounds. To thoroughly investigate potential backgrounds, we utilize an inclusive MC sample comprising 10 billion $J/\psi$ events. The inclusive MC sample includes both the production of the $J/\psi$ resonance and the continuum processes incorporated in kkmc JADACH2000260; PhysRevD.63.113009. All particle decays are modeled with evtgen LANGE2001152 using the branching fractions either taken from the Particle Data Group (PDG) PhysRevD.110.030001, when available, or otherwise estimated with lundcharm PhysRevD.62.034003. Final state radiation (FSR) from charged final state particles is incorporated using the photos package BARBERIO1991115. Specific generators using the appropriate theoretical amplitudes for $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\ell^{+}\ell^{-}$ Zhang:2012gt, $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ BESIII:2014bgm, and $\eta\rightarrow\gamma e^{+}e^{-}$ Qin:2017vkw decays were developed for this analysis.
III EVENT SELECTION
The final state of interest is studied through the decay chain $J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$, and $\eta\rightarrow\ell^{+}\ell^{-}$. Each event is required to contain at least one good photon candidate and four charged track candidates with a total charge of zero. Each charged track reconstructed from MDC hits is required to have a polar angle in the range $|\cos\theta|\leq 0.93$ and to pass the interaction point within $\pm 10$ cm along the beam direction and within $\pm 1$ cm in the plane perpendicular to the beam.
Photon candidates are reconstructed using isolated clusters of energy deposited in the EMC and must have a deposited energy larger than 25 MeV in the barrel region ($|\cos\theta|\leq 0.80$) and 50 MeV in the end-cap region ($0.86<|\cos\theta|<0.92$). The energy deposited in nearby TOF counters is included to improve the reconstruction efficiency and energy resolution. To eliminate clusters associated with charged tracks, the angle between the photon candidate and the extrapolation of any charged track to the EMC must be larger than 10 degrees. A requirement on the EMC cluster timing with respect to the event start time $(0\leq T\leq 700\,\text{ns})$ is used to suppress electronic noise and energy deposits unrelated to the event.
For each candidate, particle identification (PID) is performed using the TOF and $\text{d}E/\text{d}x$ information, and four-constraint (4C) kinematic fits are performed imposing energy and momentum conservation under the hypotheses of $J/\psi\to\gamma\pi^{+}\pi^{-}\ell^{+}\ell^{-}$. The variable, $\chi^{2}_{\rm 4C+PID}=\chi^{2}_{4\rm C}+\chi^{2}_{\rm PID}$, is used to select the best candidate in events containing more than one combination. Here $\chi^{2}_{4\rm C}$ is the $\chi^{2}$ value of the 4C kinematic fit and $\chi^{2}_{\rm PID}$ is a sum over the four charged track candidates. For each event, the hypothesis with the smallest $\chi^{2}_{\rm 4C+PID}$ is kept for further analysis.
IV Analysis of $\eta\rightarrow\mu^{+}\mu^{-}$
For the $\eta\rightarrow\mu^{+}\mu^{-}$ decay, $\chi^{2}_{\rm 4C+PID}(\pi^{+}\pi^{-}\mu^{+}\mu^{-})<40$ is required. A 4C kinematic fit under the hypothesis of $J/\psi\rightarrow\gamma 2(\pi^{+}\pi^{-})$ is also performed, and $\chi^{2}_{\rm 4C+PID}(\pi^{+}\pi^{-}\pi^{+}\pi^{-})>\chi^{2}_{4\rm C+PID}(\pi^{+}\pi^{-}\mu^{+}\mu^{-})$ is required to reject possible background events from $J/\psi\rightarrow\gamma 2(\pi^{+}\pi^{-})$. The muon counter information is not used due to the low momentum of the muons of interest.
The resulting $\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ invariant mass distribution ($M_{\pi^{+}\pi^{-}\mu^{+}\mu^{-}}$) is shown in Fig. 2. A clear $\eta^{\prime}$ peak is observed, while the small peak around 0.93 $\text{GeV}/c^{2}$ comes from the background $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ decay. The distribution of $M_{\pi^{+}\pi^{-}\mu^{+}\mu^{-}}$ versus the $\mu^{+}\mu^{-}$ invariant mass ($M_{\mu^{+}\mu^{-}}$) is displayed in Fig. 3. The cluster within the marked region corresponds to the $\eta^{\prime}\to\pi^{+}\pi^{-}\eta$ decay. Requiring $M_{\pi^{+}\pi^{-}\mu^{+}\mu^{-}}\in(0.945,0.970)~\text{GeV}/c^{2}$, an evident $\eta$ peak is seen in the $M_{\mu^{+}\mu^{-}}$ distribution, as shown in Fig. 4. Although the $\eta^{\prime}$ peak in Fig. 2 is dominated by background events from $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$, these events are primarily concentrated in the low-energy region of $M_{\mu^{+}\mu^{-}}$ and contribute very little in the mass range of interest, as shown in Fig. 4.
Fig. 2: The distribution of $M_{\pi^{+}\pi^{-}\mu^{+}\mu^{-}}$. The dots with error bars represent the data, and the red histogram is the signal MC sample. The dashed green line, dashed pink line, and shaded blue histogram represent the backgrounds from $\eta\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$, $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ and $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$, respectively.
Fig. 3: The distribution of $M_{\pi^{+}\pi^{-}\mu^{+}\mu^{-}}$ versus $M_{\mu^{+}\mu^{-}}$ for data.
To investigate background events in the $M_{\mu^{+}\mu^{-}}$ distribution, an inclusive MC sample of 10 billion $J/\psi$ events is analyzed with the same procedure. The background events are mainly from the decays $J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$, $J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ and $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$. Dedicated MC samples for these background channels are generated to estimate their contributions. We find that none of these possible background channels contributes to the $\eta$ peak, as seen in Fig. 4.
To obtain the signal yield of $\eta\rightarrow\mu^{+}\mu^{-}$, an extended unbinned maximum likelihood fit is performed to the $M_{\mu^{+}\mu^{-}}$ distribution. In the fit, the total probability density function consists of a signal component and various background contributions. The signal component is modeled with the simulated shape. The background components considered are subdivided into two classes. For the first, the shapes of the background events from $\eta^{\prime}$ decays $J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ and $J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ are taken from dedicated MC simulations. The magnitudes of these backgrounds are fixed at $15.0\pm 3.0$ and $0.8\pm 0.1$, respectively, normalized using the branching fractions of $J/\psi\rightarrow\gamma\eta^{\prime}$ and their subsequent decays. The second class is the $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ component, also described with the MC-simulated shape, with a magnitude left as a free parameter in the fit.
[…]
$\mathcal{B}(\eta\rightarrow\mu^{+}\mu^{-})$ $\,=\,\frac{N_{\rm obs}}{N_{J/\psi}\,{\mathcal{B}}_{\rm int}\,\varepsilon_{\eta\rightarrow\mu^{+}\mu^{-}}}$ (1)
[…]
where $N_{\rm obs}$ is the signal yield, $N_{J/\psi}$ is the total number of $J/\psi$ events, $\mathcal{B}_{\rm int}\equiv\mathcal{B}(J/\psi\rightarrow\gamma\eta^{\prime})\,\mathcal{B}(\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta)$ is taken from the PDG PhysRevD.110.030001, $\varepsilon_{\eta\rightarrow\mu^{+}\mu^{-}}$ is the detection efficiency, and the uncertainty is statistical only.
Fig. 4: The fit to the $M_{\mu^{+}\mu^{-}}$ distribution. The dots with error bars represent the data. The blue line represents the total fit result, and the red line represents the signal component. The dashed green line, shaded pink shape, and dot-dash blue line are the contributions from the decays $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$, $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ and $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$, respectively.
V Analysis of $\eta\rightarrow e^{+}e^{-}$
To select $\eta\to e^{+}e^{-}$ events, the $\chi^{2}_{4\rm C+PID}(\pi^{+}\pi^{-}e^{+}e^{-})$ is required to be less than 40. The $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass ($M_{\pi^{+}\pi^{-}e^{+}e^{-}}$) is shown in Fig. 5(a). No evident $\eta^{\prime}$ signal is observed. We then select the $\eta^{\prime}$ mass window to be within $(0.945,0.97)\,\text{GeV}/c^{2}$. Fig. 5(b) shows the $e^{+}e^{-}$ invariant mass ($M_{e^{+}e^{-}}$) distribution for candidates passing all selection criteria. No events are observed in the $\eta$ mass region.
A background study based on the inclusive MC sample indicates that the dominant background originates from $J/\psi\rightarrow\gamma\eta^{\prime}$ with the subsequent decays $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$, $\eta\rightarrow\gamma e^{+}e^{-}$ and $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$. Dedicated MC samples are generated to evaluate contributions from these channels. The background yields are $0.5\pm 0.3$ and $1.0\pm 0.2$ events, respectively.
The signal region is defined as the interval within 3.5 standard deviations. The standard deviation represents the mass resolution obtained from signal MC events, corresponding to a mass interval of $(0.534,0.560)\,\text{GeV}/c^{2}$. No events are observed in this interval, $N_{\rm sig}=0$. The background contribution, $N_{\rm bkg}$, in the signal region is estimated using events in sideband regions. Symmetric sideband regions, each with the same width as the signal region, are defined as $(0.508,0.534)\,\text{GeV}/c^{2}$ and $(0.560,0.586)\,\text{GeV}/c^{2}$, giving $N_{\rm bkg}=1.5$. With the dedicated MC sample, the detection efficiency is determined to be $(27.48\pm 0.04)\%$, where the uncertainty is statistical only.
The upper limit on the signal yield, $N_{\rm UL}$, is estimated at a CL of 90% using the method from Ref. Rolke:2000ij, which is suitable for small signals in the presence of background. Given $N_{\rm sig}$ and $N_{\rm bkg}$, $N_{\rm UL}$ of $\eta\to e^{+}e^{-}$ is estimated to be 1.3. Accounting for efficiency and systematic uncertainty $\delta_{\rm sys}$ (see Section VI below for details), the upper limit on the branching fraction of $\eta\to e^{+}e^{-}$ is determined as
$\mathcal{B}(\eta\rightarrow e^{+}e^{-})$ $\,<\,\frac{N_{\rm UL}}{N_{J/\psi}\,\mathcal{B}_{\rm int}\,\varepsilon_{\eta\to e^{+}e^{-}}\,(1-\delta_{\rm sys})}$ (2)
[…]
Fig. 5: The distributions of (a) $M_{\pi^{+}\pi^{-}e^{+}e^{-}}$ and (b) $M_{e^{+}e^{-}}$. Dots with error bars represent the data, the red histogram is the signal MC shape. Two blue arrows represent the signal region.
VI Systematic Uncertainties
The sources of systematic uncertainties include the efficiency differences between the data the and MC simulation for the MDC tracking, the PID, the photon detection, and the kinematic fit, as well as the fitting procedure and the total number of $J/\psi$ events. Their effects on the branching fractions are discussed below.
  • Total number of $J/\psi$ events:
    The total number of $J/\psi$ events, (10087 $\pm$ 44) $\times$ 106 BESIII:2021cxx, is determined from an analysis of inclusive $J/\psi$ hadronic events, and the uncertainty, 0.4%, is taken as the systematic uncertainty.
  • Uncertainties of $\mathcal{B}(J/\psi\rightarrow\gamma\eta^{\prime})$ and $\mathcal{B}(\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta)$:
    The branching fractions of $J/\psi\rightarrow\gamma\eta^{\prime}$ and $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$ are taken from the PDG PhysRevD.110.030001. Their uncertainties are 1.3% and 1.2%, respectively.
  • MDC tracking:
    The single charged pion track efficiency has been studied with respect to the transverse momentum and polar angle $\theta$ in the MDC by means of a high-purity sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0}$ liuStudyTrackingEfficiency2025 obtained by using one pion to tag the presence of the other. Two-dimensional tracking efficiencies as a function of the transverse momentum and $\cos\theta$ are evaluated for both the data and MC simulation. To correct the MC efficiencies to those of the data, the corresponding two-dimensional correction factors and their uncertainties are derived. To estimate its impact on the systematic uncertainty, the MC simulation events are then weighted to determine the detection efficiency. Since the mass and the behavior of low-momentum muons are similar to those of pions, we weight the muon tracks by assuming them to be pions.
    In the case of electrons, the clean samples of $e^{+}e^{-}\rightarrow e^{+}e^{-}\gamma$ at the $J/\psi$ meson mass and $J/\psi\rightarrow e^{+}e^{-}\gamma_{\rm FSR}$ are used to evaluate the MC-data tracking difference as a function of momentum and polar angle, which are then used to weight the MC simulated events. The changes in the detection efficiencies with and without correction, 0.8% for $\eta\rightarrow\mu^{+}\mu^{-}$ and 2.1% for $\eta\rightarrow e^{+}e^{-}$, are considered as the systematic uncertainties.
  • PID and 4C kinematic fit:
    In case of $\eta\rightarrow\mu^{+}\mu^{-}$, the uncertainties from PID and the kinematic fit are evaluated separately. To estimate the pion PID efficiency uncertainty, we have studied the pion PID efficiency using the control samples of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0}$. The results show that the pion PID efficiency for data agrees well with that of the MC simulation. The minor MC-data discrepancy as a function of the momentum and the polar angle is used to estimate the systematic uncertainty with the same approach as for the tracking efficiency. Due to a lack of a specific control sample for low-momentum muons, the PID efficiency correction is performed with the correction factor obtained from pions. The uncertainty associated with the 4C kinematic fit is studied by correcting the track helix parameters PhysRevD.87.012002 to reduce the difference between data and MC simulation. The differences between the efficiencies with and without corrections, summarized in Table 1, are taken as the systematic uncertainties from PID and 4C kinematic fit.
    For $\eta\rightarrow e^{+}e^{-}$, we use a clean control sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0},\pi^{0}\rightarrow\gamma e^{+}e^{-}$ to study the uncertainty. This control sample has the same topology as the decay of interest. Using the same charged tracks and photon selection criteria, we first select the candidate events with a very loose requirement on $\chi^{2}_{4\rm C+PID}$. Then, we further tighten the selection by requiring $\chi^{2}_{4\rm C+PID}<40$. The efficiency difference between data and MC simulation, 4.6%, is taken as the systematic uncertainty.
    Table 1: Relative systematic uncertainties (in %) in the branching fraction measurements.
    Source $\eta\rightarrow\mu^{+}\mu^{-}$ $\eta\rightarrow e^{+}e^{-}$
    Number of $J/\psi$ events $0.4$ $0.4$
    $\mathcal{B}(J/\psi\rightarrow\gamma\eta^{\prime})$ $1.3$ $1.3$
    $\mathcal{B}(\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta)$ $1.2$ $1.2$
    MDC tracking $0.8$ $2.1$
    PID and kinematic fit 2.7 $4.6$
    Photon detection $0.5$ $0.5$
    Signal shape $0.1$ -
    Background shape $0.2$ -
    Total 3.4 5.6
  • Photon detection efficiency:
    The systematic uncertainty from the reconstruction of photons has been extensively studied in the processes $e^{+}e^{-}\rightarrow\gamma\mu^{+}\mu^{-}$. The relative difference in efficiency between data and MC simulation, up to 0.5%, is assigned as the systematic uncertainty.
  • Signal shape:
    The uncertainties due to the signal MC shape is estimated by performing an alternative fit in which the signal is modeled by the MC shape convolved with a single Gaussian function. It is found that the difference of the determined number of signal events, 0.1%, which is taken as the systematic uncertainty.
  • Background shape:
    In the fit to $M_{\mu^{+}\mu^{-}}$, the numbers of background events from $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ and $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ are fixed at the normalized values according to the branching fractions from the PDG PhysRevD.110.030001. To estimate the effect of the uncertainty of the used branching fractions, the numbers of events for these two backgrounds are varied randomly within the uncertainties of the branching fractions. Using these randomly scaled parameters, we perform a series of fits to the $M_{\mu+\mu^{-}}$ distribution. The difference of the determined number of signal events, 0.2%, is taken as the systematic uncertainty.
  • Fit range:
[…]
Table 1 summarizes all contributions to the systematic uncertainties on the branching fraction measurements. In each case, the total systematic uncertainty is obtained by adding the individual contributions in quadrature.
VII Summary
With a sample of (10087 $\pm$ 44) $\times$ 106  BESIII:2021cxx $J/\psi$ events collected with the BESIII detector, we perform a study of $\eta\to\ell^{+}\ell^{-}$ tagging with $\eta^{\prime}\to\pi^{+}\pi^{-}\eta$ in $J/\psi$ radiative decays. The decay $\eta\rightarrow\mu^{+}\mu^{-}$ is observed with significance of $9.8\sigma$, and its decay branching fraction is measured to be $(5.8\pm 1.0_{\rm stat}\pm 0.2_{\rm syst})\times 10^{-6}$. This is consistent with theoretical predictions Petri:2010ea and the $(5.8\pm 0.8)\times 10^{-6}$ PDG value PhysRevD.110.030001. No signal from the $\eta\rightarrow e^{+}e^{-}$ is observed, and the upper limit on its decay branching fraction at 90% CL is set at $2.2\times 10^{-7}$. The result improves the upper limit of $7\times 10^{-7}$ $(CL=90\%)$ from the SND experiment SND:2018egu, providing a stricter experimental constraint on new physics.
In the future, new experimental facilities, such as the super $\tau$-charm facility (STCF) Achasov_2023, the JLab Eta Factory (JEF) Somov:2024jiy, REDTOP Gatto:2019dhj and the $\eta$ factory of HIAF An:2025lws experiments, will provide larger $\eta$ samples. They are expected to conduct precise studies of the $\eta\rightarrow e^{+}e^{-}$ decay, and thereby deeply explore the internal structure of the $\eta$ meson, and search for new physics phenomena beyond the SM.
VIII Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017,12475089 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; ERC under Contract No. 758462; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; STFC (United Kingdom); Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374